# Soteria: Language-Specific Functional Parameter Steering for Multilingual Safety Alignment

### v1

@@ Abstract
Ensuring consistent safety across multiple languages remains a significant challenge for large language models (LLMs). We introduce Soteria, a lightweight yet powerful strategy that locates and minimally adjusts the “functional heads” most responsible for harmful content generation in each language. By altering only a fraction of parameters, Soteria drastically reduces policy violations without sacrificing overall model performance, even in low-resource settings. To rigorously evaluate our approach, we also present XThreatBench, a specialized multilingual dataset capturing fine-grained harmful behaviors drawn from real policy guidelines. Experiments with leading open-source LLMs (e.g., Llama, Qwen, Mistral) show that Soteria consistently improves safety metrics across high-, mid-, and low-resource languages. These findings highlight a promising path toward scalable, linguistically attuned, and ethically aligned LLMs worldwide. The source code and dataset are available at: https://github.com/NeuralSentinel/Soteria
[…]
[ topline=false, bottomline=false, skipabove=skipbelow=leftline=true, rightline=true, linecolor=NavyBlue, linewidth=2pt, innertopmargin=5pt, innerbottommargin=5pt, innerrightmargin=5pt, innerleftmargin=5pt, backgroundcolor=gray!10, roundcorner=10pt ]stylishframe
Soteria: Language-Specific Functional Parameter Steering for Multilingual Safety Alignment
Somnath Banerjee, Sayan Layek, Pratyush Chatterjee, Animesh Mukherjee, Rima Hazra Indian Institute of Technology Kharagpur INSAIT, Sofia University “St. Kliment Ohridski”
@@ 1 Introduction
LLMs such as GPT-4o OpenAI et al. (2024), Claude Anthropic , and Llama Touvron et al. (2023) have revolutionized the AI landscape by delivering impressive performance across tasks ranging from text generation to question answering. These breakthroughs stem from extensive pre-training on large, diverse corpora Zhou et al. (2023); Kamalloo et al. (2023); Nguyen et al. (2024a). Yet, much of the early research on LLMs’ multilingual capabilities relied on translating English queries into non-English, a strategy that obscures genuine multilingual performance Zhao et al. (2024). Although newer LLMs feature advanced tokenizers that handle non-English inputs more effectively, key safety measures, including red teaming and content filtering remain predominantly English centric Zhang et al. (2023); Gurgurov et al. (2024).
As a result, non-English use cases are comparatively under-protected, and especially smaller-parameter models (e.g., 8B or 7B) often implemented in low-resource settings are at greater risk of generating harmful or culturally insensitive outputs Banerjee et al. (2025). Moreover, prior work on safety mechanisms has focused mainly on English, overlooking the nuances and needs of broader linguistic communities Banerjee et al. (2024b); Hazra et al. (2024a). In this context, it becomes clear that robust, multilingual safety protocols are essential to protect users and maintain linguistic sensitivity across the globe Wang et al. (2024); Lu and Koehn (2024).
A major obstacle to robust multilingual safety lies in the limitations of early tokenizers Petrov et al. (2023); Hong et al. (2024), which were not designed properly to capture the rich morphological and script diversity in global languages Ali et al. (2024). As a result, LLMs built on these tokenizers struggle to generate linguistically relevant and accurate outputs in non-English settings, undermining the effectiveness of any safety measures. While newer models incorporate more sophisticated multilingual tokenizers [^1] https://huggingface.co/blog/llama31, prior efforts largely treated multilingual support as an afterthought added later via fine-tuning rather than integrated as a core capability Richburg and Carpuat (2024). This approach often relies on “bridging strategies,” such as translating queries into English before applying moderation filters, a practice that can distort content classification Bang et al. (2023); Lai et al. (2024). Even extensive fine-tuning typically fails to address deeper, English-dominant architectural constraints, especially for languages with multiple scripts or highly complex morphology. Moreover, creating large-scale multilingual datasets for each fine-tuning cycle is prohibitively expensive and time-intensive Yu et al. (2022). Although scaling up to larger-parameter models can bolster multilingual proficiency, such approaches may be infeasible in low-resource or time-sensitive contexts Nguyen et al. (2024b); Chelombitko et al. (2024).
[…]
  • To the best of our knowledge, we are the first to introduce a multilingual parameter-efficient safety mechanism – Soteria – that modifies only about $\sim$3of the model’s language-specific “functional heads,” effectively reducing harmful outputs without compromising overall performance.
@@ 2 Related work
Safety alignment: Efforts to ensure LLM safety focus on mitigating adversarial prompts Xie et al. (2018), designing robust filtering Xiao et al. (2024), and maintaining dynamic oversight Kenton et al. (2024); Wang et al. (2024). Early studies Yao et al. (2024) expose key vulnerabilities and propose ethical risk frameworks. Subsequent work Sachdeva et al. (2025); Banerjee et al. (2024a) reveals how subtle prompt manipulations can evade safeguards, prompting research into attack strategies Wolf et al. (2024) and defenses like RAIN Li et al. (2023). Others emphasize dynamic monitoring Bhardwaj et al. (2024) and adaptive safety mechanisms, including safety arithmetic Hazra et al. (2024b) for test-time alignment and SafeInfer Banerjee et al. (2024b), SafeDecoding Xu et al. (2024) for decoding-time alignment.
@@ 3.1 Preliminaries
We define an autoregressive LLM as $\mathcal{M}$, which comprises multiple transformer decoder layers, denoted by $\mathcal{L}$. Each transformer decoder layer consists of two fundamental modules – multi-head attention ($MHA$) and feed-forward network ($FFN$). The outputs of $MHA$ and $FFN$ modules in layer $l\in\mathcal{L}$ are denoted by $atn^{l}$ and $mlp^{l}$, respectively. The hidden state of a transformer decoder layer $l$ is denoted by $ht_{l}$. The hidden state $ht_{l}$ is computed as shown in Equation 1 where $ht_{l-1}$ represents the hidden state from the previous layer $l-1$.
@@
$\vspace{-0.2cm}ht_{l}=ht_{l-1}+mlp^{l}+atn^{l}$ (1)
Mathematically, the output $atn^{l}$ of $MHA$ module is further obtained using Equation 2 in which each attention head is represented as $h_{i}^{l}$ where $i\in\mathcal{I}$ denotes the $i^{th}$ attention head and $|\mathcal{I}|$ denotes the number of heads in each layer $l$. $W^{O}_{l}\in\mathbb{R}^{|\mathcal{I}|\cdot d_{k}\times d_{m}}$ projects ($O$ Projection) the concatenated heads to the model dimension whereby the head $h_{i}^{l}$ has a dimension of $d_{k}$ and the hidden dimension of the model is $d_{m}$. Each head $h_{i}^{l}$ is derived as given in Equation 3 in which $W_{i}^{Q}$, $W_{i}^{K}$ and $W_{i}^{V}$ denote the learned weight matrices for the query $Q$, key $K$, and values $V$ of the $i^{th}$ head.
$\vspace{-0.2cm}atn_{l}=\text{concat}(h_{1}^{l},\dots,h_{\mathcal{I}}^{l})\cdot W^{O}_{l}$ (2)
@@
In this work, similar to Todd et al. (2024), we adopt the attention definition proposed by Elhage et al. (2021) rather than the one introduced in Vaswani et al. (2017). The study in  Elhage et al. (2021) highlights that the formulation in  Vaswani et al. (2017) can be interpreted as decomposing weight matrix $W^{O}_{l}$ into a block form $[W_{l1}^{O}\;W_{l2}^{O}\;\dots\;W_{l\mathcal{I}}^{O}]$, allowing $h_{i}^{l}$ to be directly projected into residual stream space. Each block $W_{li}^{O}\in\mathbb{R}^{d_{k}\times d_{m}}$ determines how information from $h_{i}^{l}$ is transformed into the final model dimension. We use the output $atn_{i}^{l}$ corresponding to $i^{th}$ head as written in Equation 4.
$\vspace{-0.2cm}atn_{i}^{l}=h_{i}^{l}\cdot W_{li}^{O}\in\mathbb{R}^{d_{m}}$ (4)
@@ 3.2 Our framework
In our framework, we first identify important attention heads (i.e., $atn_{i}^{l}$ for the $i^{th}$ head) and subsequently remove the harm direction from the target model.
@@
$\vspace{-0.2cm}\hat{atn_{i}^{l}}_{t}=\frac{1}{|\mathscr{P}_{t}|}\sum_{p_{k}^{t}\in\mathscr{P}_{t}}atn_{i}^{l}(p_{k}^{t})$ (5)
In parallel, we have a corrupted prompt $\hat{p}_{i}^{k}$ where the responses are shuffled $\hat{p}_{i}^{k}=\left[(q_{k_{1}},\hat{r}_{k_{1}}),\cdots,(q_{k_{K}},\hat{r}_{k_{K}}),q_{k_{Q}}\right]$. Next, we pass the corrupted prompt $\hat{p}_{k}^{t}$ through the language model $\mathcal{L}$ and replace a specific attention head activation $atn_{i}^{l}(\hat{p}_{k}^{t})$ with the actual mean task conditioned activation $\hat{atn_{i}^{l}}_{t}$. We attempt to understand how much the actual task conditioned activation can help to predict the correct answer. Further we measure the causal indirect effect (CIE) toward recovering the correct answer $r_{k_{Q}}$ as shown in Equation 6.
@@
Removal of harm direction: According to Equation 4, each block $W_{li}^{O}$ determines the transformation of information from $h_{i}^{l}$ to the output $atn^{l}_{i}$. Given an important attention $atn^{l}_{i}$, we consider the associated block $W_{li}^{O}$ for harm direction removal. We focus solely on the $O$-projection weight, avoiding unnecessary changes to other layer weights, which could compromise the model’s broader capabilities. Following Hazra et al. (2024a) we compute the harm vector $H_{v}$ by taking the element-wise difference between the $\mathcal{M}_{H}$ and $\mathcal{M}$. Further, we keep only those parameters of $H_{v}$ as per selected blocks ($W_{li}^{O}$ for $i^{th}$ head) of the $W^{O}_{l}$ and make the other parameters zero. The harm vector with retained parameters is denoted by $\hat{H}_{v}$. The safe model $\hat{\mathcal{M}}$ is expressed as follows.
^$\vspace{-0.2cm}\boxed{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\hat{\mathcal{M}}}=\mathcal{M}-\lambda*{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\hat{H}_{v}}}$ (7)
@@ 4 Language and dataset
Datasets: We assess Soteria using two established datasets, MultiJail Deng et al. (2024b) and XSafety Wang et al. (2024). In addition, we introduce a new multilingual safety dataset XThreatBench, constructed based on the policy violations outlined by Meta Qi et al. (2023a). A detailed description of each dataset follows. We include the dataset details of XSafety and the corresponding experimental results in the Appendix A due to space constraints.
@@
XThreatBench: We propose a multilingual safety benchmark on general harm designed to rigorously evaluate LLM vulnerabilities across 10 high-risk categories, including sexual content, child exploitation, economic fraud, hate speech, illegal activities, cyber threats, physical harm, political manipulation, privacy violations, and deception. XThreatBench features 3,000 adversarial prompts across 12 languages ensuring native linguistic authenticity with generic harm. Each category contains 25 instances, systematically crafted to test LLM safety mechanisms against general attack scenarios, evasive jailbreak tactics, and multilingual exploits. To maximize adversarial quality, we implemented a three-stage verification process, combining partial human intervention, GPT-4 based filtering, and Perspective API based toxicity scoring (retaining only queries with a toxicity score of 0.75+). Unlike existing benchmarks, XThreatBench adheres to safety policies while offering a robust, multilingual evaluation framework for assessing LLM safety in high-risk, adversarial environments.
@@ 5.1 Language models
We use four open-weight language models to evaluate our proposed safety mechanism. We consider instruction-tuned versions of llama3.1-8b-instruct Grattafioriet et al. (2024), Qwen2-7B-Instruct Yang et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), Phi-3.5-mini-instruct Abdin et al. (2024). We select these models because they are widely used instruction-tuned LLMs with diverse multilingual capabilities. They perform well in high-resource languages such as English and Chinese, while their support for medium- and low-resource languages varies. This selection allows us to evaluate safety mechanisms across different linguistic distributions and understand their effectiveness in multilingual settings.
5.2 Language specific tasks used
Following the tasks outlined in Todd et al. (2024), we use five datasets: (1) agnews, a text classification dataset of news headlines and early article sentences labeled as business, science/technology, sports, or world; (2) antonym, where each input word requires generating its opposite; (3) country-capital, comprising 197 country-capital pairs and requiring the capital city given a country; (4) country-currency, listing countries alongside their currencies; and (5) sentiment, which contains 1167 movie review sentences labeled as either positive or negative. Further, we translate these task datasets into specific languages and identify the attention heads.
5.3 Evaluation metrics
Following previous studies Chao et al. (2024); Qi et al. (2023b), we assess the safety standards of language models using the attack success rate (ASR). ASR measures the proportion of harmful responses generated by the model out of the total number of harmful queries and is defined as follows.
ASR=# responses not aligned with Safety# input queries to $\text{ASR}=\frac{\text{\# responses not aligned with {Safety}}}{\text{\# inputqueries to LLM}}$
@@ 5.3 Evaluation metrics
5.3.1 GPT-4o as a safety evaluator
To reliably and efficiently assess a model’s safety performance, we employ GPT-4o as the annotator, leveraging its strong alignment with human judgments on harmful responses Qi et al. (2023b); Banerjee et al. (2024b); Hazra et al. (2024a). In particular, we adopt the evaluation prompt [^5] Given the following text, analyze and assess its ethical implications. Determine whether the content adheres to ethical standards or if there are elements that raise ethical concerns. proposed by Banerjee et al. (2024b). Cases in which GPT-4o declines to provide annotations due to content filtering are excluded from our calculations. A lower ASR score indicates a safer model.

### v2

@@ Abstract
Ensuring consistent safety across multiple languages remains a significant challenge for large language models (LLMs). We introduce Soteria, a lightweight yet powerful strategy that locates and minimally adjusts the “functional heads” most responsible for harmful content generation in each language. By altering only a fraction of parameters, Soteria drastically reduces policy violations without sacrificing overall model performance, even in low-resource settings. To rigorously evaluate our approach, we also present XThreatBench, a specialized multilingual dataset capturing fine-grained harmful behaviors drawn from real policy guidelines. Experiments with leading open-source LLMs (e.g., Llama, Qwen, Mistral) show that Soteria consistently improves safety metrics across high-, mid-, and low-resource languages. These findings highlight a promising path toward scalable, linguistically attuned, and ethically aligned LLMs worldwide. We release the source codes at: https://github.com/neuralsentinel/soteria.
Soteria: Language-Specific Functional Parameter Steering
for Multilingual Safety Alignment
Somnath Banerjee, Sayan Layek, Pratyush Chatterjee, Animesh Mukherjee, Rima Hazra Indian Institute of Technology Kharagpur, India Eindhoven University of Technology, Netherlands
@@ 1 Introduction
A major obstacle to robust multilingual safety lies in the limitations of early tokenizers Petrov et al. (2023); Hong et al. (2024), which were not designed properly to capture the rich morphological and script diversity in global languages Ali et al. (2024). As a result, LLMs built on these tokenizers struggle to generate linguistically relevant and accurate outputs in non-English settings, undermining the effectiveness of any safety measures. While newer models incorporate more sophisticated multilingual tokenizers [^1] https://huggingface.co/blog/llama31, prior efforts largely treated multilingual support as an afterthought added later via fine-tuning rather than integrated as a core capability Richburg and Carpuat (2024). This approach often relies on “bridging strategies,” such as translating queries into English before applying moderation filters, a practice that can distort content classification Bang et al. (2023); Lai et al. (2024). Even extensive fine-tuning typically fails to address deeper, English-dominant architectural constraints, especially for languages with multiple scripts or highly complex morphology. Moreover, creating large-scale multilingual datasets for each fine-tuning cycle is prohibitively expensive and time-intensive Yu et al. (2022). Although scaling up to larger-parameter models can bolster multilingual proficiency, such approaches may be infeasible in low-resource or time-sensitive contexts Nguyen et al. (2024); Chelombitko et al. (2024).
Building on these insights, we focus on recently introduced models, which offer improved multilingual capability. We curate a specialized dataset XThreatBench of prohibited categories, derived from Meta’s content guidelines to identify safety concerns more accurately. Using this dataset, we propose Soteria, a novel strategy for safe multilingual generation that locates language-specific “functional heads” and selectively tunes only about $\sim$3% of the model parameters. By redirecting these heads away from harmful outputs, Soteria effectively suppresses toxic or policy-violating responses without degrading overall model performance. Through this precise calibration of multilingual fluency and safety, we demonstrate that LLMs can be both linguistically adaptive and ethically grounded. Our contributions are as follows.
  • To the best of our knowledge, we are the first to introduce a multilingual parameter-efficient safety mechanism – Soteria – that modifies only about $\sim$3% of the model’s language-specific “functional heads,” effectively reducing harmful outputs without compromising overall performance.
@@ 2 Related work
Safety alignment: Efforts to ensure LLM safety focus on mitigating adversarial prompts Xie et al. (2018), designing robust filtering Xiao et al. (2024), and maintaining dynamic oversight Kenton et al. (2024); Wang et al. (2024). Early studies Yao et al. (2024) expose key vulnerabilities and propose ethical risk frameworks. Subsequent work Sachdeva et al. (2025); Banerjee et al. (2024a) reveals how subtle prompt manipulations can evade safeguards, prompting research into attack strategies Wolf et al. (2024) and defenses like RAIN Li et al. (2023). Others emphasize dynamic monitoring Bhardwaj et al. (2024) and adaptive safety mechanisms, including safety arithmetic Hazra et al. (2024a) for test-time alignment and SafeInfer Banerjee et al. (2024b), SafeDecoding Xu et al. (2024) for decoding-time alignment.
@@ 3.1 Preliminaries
We define an autoregressive LLM as $\mathcal{M}$, which comprises multiple transformer decoder layers, denoted by $\mathcal{L}$. Each transformer decoder layer consists of two fundamental modules – multi-head attention ($MHA$) and feed-forward network ($FFN$). The outputs of $MHA$ and $FFN$ modules in layer $l\in\mathcal{L}$ are denoted by $atn^{l}$ and $mlp^{l}$, respectively. The hidden state of a transformer decoder layer $l$ is denoted by $ht_{l}$. The hidden state $ht_{l}$ is computed as shown in Equation 1 where $ht_{l-1}$ represents the hidden state from the previous layer l1l-1.
@@
$ht_{l}=ht_{l-1}+mlp^{l}+atn^{l}$ (1)
Mathematically, the output $atn^{l}$ of $MHA$ module is further obtained using Equation 2 in which each attention head is represented as $h_{i}^{l}$ where $i\in\mathcal{I}$ denotes the $i^{th}$ attention head and $|\mathcal{I}|$ denotes the number of heads in each layer $l$. $W^{O}_{l}\in\mathbb{R}^{|\mathcal{I}|\cdot d_{k}\times d_{m}}$ projects ($O$ - Projection) the concatenated heads to the model dimension whereby the head $h_{i}^{l}$ has a dimension of $d_{k}$ and the hidden dimension of the model is $d_{m}$. Each head $h_{i}^{l}$ is derived as given in Equation 3 in which $W_{i}^{Q}$, $W_{i}^{K}$ and $W_{i}^{V}$ denote the learned weight matrices for the query $Q$, key $K$, and values $V$ of the $i^{th}$ head.
$atn_{l}=\text{concat}(h_{1}^{l},\dots,h_{\mathcal{I}}^{l})\cdot W^{O}_{l}$ (2)
@@
In this work, similar to Todd et al. (2024), we adopt the attention definition proposed by Elhage et al. (2021) rather than the one introduced in Vaswani et al. (2017). The study in  Elhage et al. (2021) highlights that the formulation in  Vaswani et al. (2017) can be interpreted as decomposing weight matrix $W^{O}_{l}$ into a block form $W_{l1}^{O}\;W_{l2}^{O}\;\dots\;W_{l\mathcal{I}}^{O}]$, allowing $h_{i}^{l}$ to be directly projected into residual stream space. Each block $W_{li}^{O}\in\mathbb{R}^{d_{k}\times d_{m}}$ determines how information from $h_{i}^{l}$ is transformed into the final model dimension. We use the output $atn_{i}^{l}$ corresponding to $i^{th}$ head as written in Equation 4.
$atn_{i}^{l}=h_{i}^{l}\cdot W_{li}^{O}\in\mathbb{R}^{d_{m}}$ (4)
@@
3.2 Why modify attention heads?
Decoder-only transformer architectures compute attention scores to capture pairwise interactions between tokens in the input sequence via self-attention. This mechanism allows each token to condition directly on its prior context. As such, attention heads naturally mediate how past tokens influence the generation of the next token. Consequently, attention heads in LLM decoders are ideal intervention points for fine-grained control over model behavior.
Recent work has established that a small subset of attention heads disproportionately contribute to solving specific tasks Todd et al. (2024); Zhou et al. (2025); Banerjee et al. (2024b). Notably, Zhou et al. (2025) empirically showed that the top task-relevant attention heads also correlate with heads that are safety-critical. This motivates our design to target only such functional heads, rather than the entire model.
@@ 3.3 Our framework
In our framework (see Figure 1), we first identify important attention heads (i.e., $atn_{i}^{l}$ for the $i^{th}$ head) and subsequently remove the harm direction from the target model.
@@
$\vskip-5.69046pt\hat{atn_{i}^{l}}_{t}=\frac{1}{|\mathscr{P}_{t}|}\sum_{p_{k}^{t}\in\mathscr{P}_{t}}atn_{i}^{l}(p_{k}^{t})$ (5)
In parallel, we have a corrupted prompt $\hat{p}_{i}^{k}$ (see Appendix for examples) where the responses are shuffled $\hat{p}_{i}^{k}=\left[(q_{k_{1}},\hat{r}_{k_{1}}),\cdots,(q_{k_{K}},\hat{r}_{k_{K}}),q_{k_{Q}}\right]$. Next, we pass the corrupted prompt $\hat{p}_{k}^{t}$ through the language model $\mathcal{L}$ and replace a specific attention head activation $atn_{i}^{l}(\hat{p}_{k}^{t})$ with the actual mean task conditioned activation $\hat{atn_{i}^{l}}_{t}$. We attempt to understand how much the actual task conditioned activation can help to predict the correct answer. Further we measure the causal indirect effect (CIE) toward recovering the correct answer $r_{k_{Q}}$ as shown in Equation 6.
@@
Removal of harm direction: According to Equation 4, each block $W_{li}^{O}$ determines the transformation of information from $h_{i}^{l}$ to the output $atn^{l}_{i}$. Given an important attention $atn^{l}_{i}$, we consider the associated block $W_{li}^{O}$ for harm direction removal. We focus solely on the $O$-projection weight, avoiding unnecessary changes to other layer weights, which could compromise the model’s broader capabilities. Following Hazra et al. (2024b) we compute the harm vector $H_{v}$ by taking the element-wise difference between the $\mathcal{M}_{H}$ and $\mathcal{M}$. Further, we keep only those parameters of $H_{v}$ as per selected blocks ($W_{li}^{O}$ for $i^{th}$ head) of the $W^{O}_{l}$ and make the other parameters zero. The harm vector with retained parameters is denoted by $\hat{H}_{v}$. The safe model $\hat{\mathcal{M}}$ is expressed as follows.
[…]
@@ 4 Language and dataset
Datasets: We assess Soteria using two established datasets, MultiJail Deng et al. (2024b) and XSafety Wang et al. (2024). In addition, we introduce a new multilingual safety dataset XThreatBench, constructed based on the policy violations outlined by Meta Qi et al. (2023a). A detailed description of each dataset follows. We include the dataset details of XSafety and the corresponding experimental results in the Appendix E due to space constraints.
@@
XThreatBench: To comprehensively evaluate multilingual safety vulnerabilities in LLMs, we introduce XThreatBench, a novel benchmark of harmful prompts grounded in real-world moderation policies. Unlike prior resources that rely on direct translations of English queries, XThreatBench is systematically constructed to ensure policy alignment, adversarial robustness, and linguistic diversity across 12 languages.
Step 1: Category derivation and prompt generation. To construct XThreatBench, we systematically consider high-risk categories outlined in Meta’s policy documents [^3] https://transparency.meta.com/en-gb/policies/ and https://about.meta.com/actions/safety/topics/safety-basics/policies/. We define 10 core categories that frequently appear in safety evaluations: sexual content, child sexual exploitation, hate speech, violence and physical harm, cybersecurity and malware, terrorism and extremism, privacy violations and doxxing, political misinformation and manipulation, deceptive behavior, and economic scams and financial harm. Each of these parent categories are further refined into granular subcategories for high-resolution threat modelling. For each subcategory, we prompt an unsafe LLM (undisclosed to avoid misuse) to generate English prompts reflecting policy-violating behaviour. These prompts serve as candidates for the harmful dataset pool.
Step 2: Filtering via GPT-4o. The generated prompts are filtered using GPT-4o to assess whether they reflect harmful intent. GPT-4o served as a first-stage semantic verifier, and we retain only the prompts it categorized as harmful. This step ensures the standards of a high-quality safety judgment scheme and helps filter out noise or benign queries.
Step 3: Toxicity scoring using Perspective API [^4] https://perspectiveapi.com/. The filtered prompts are then passed through the Perspective API to assign toxicity scores in the range $[0,1]$. We retain only those prompts with a toxicity score exceeding 0.7. This ensurs that the final dataset consists of high-confidence harmful examples only.
Step 4: Multilingual expansion. The resulting high-toxicity prompts are translated into 12 target languages using the Google Translator API. These languages span a range of typological and resource diversity, including high-resource (English, Spanish, Chinese, French, German), mid-resource (Hindi, Arabic, Bulgarian, Thai), and low-resource (Bengali, Tamil, Telugu) languages. While automatic translation iss used across the board, we manually verify a subset of queries in Bengali, Hindi, Tamil and Telugu. Given the strong annotation agreement and shared filtering pipeline, we assume similar semantic fidelity for other languages.
Dataset composition. XThreatBench contains 3,000 harmful prompts across 12 languages and 10 harm categories (see Figure 10 for examples). Each prompt includes metadata such as language, category, subcategory, GPT-4 harm judgment, and Perspective API score. The dataset is designed to facilitate cross-lingual safety evaluation under general-purpose, adversarial conditions, enabling model probing for both aligned and evasive threat scenarios.
Ethical safeguards. All prompts are synthetic and derived from publicly available moderation categories. No private or user-derived data is included. The dataset is intended exclusively for research in safety alignment, multilingual robustness, and adversarial evaluation, and adheres to established ethical standards for LLM auditing.
@@ 5 Experimental setup
Language models: We use four open-weight language models to evaluate our proposed safety mechanism. We consider instruction-tuned versions of llama3.1-8b-instruct Grattafioriet et al. (2024), Qwen2-7B-Instruct Yang et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), Phi-3.5-mini-instruct Abdin et al. (2024). We select these models because they are widely used instruction-tuned LLMs with diverse multilingual capabilities. They perform well in high-resource languages such as English and Chinese, while their support for medium- and low-resource languages varies. This selection allows us to evaluate safety mechanisms across different linguistic distributions and understand their effectiveness in multilingual settings.
Language specific tasks used: Following the tasks outlined in Todd et al. (2024), we use five datasets: (1) agnews, a text classification dataset of news headlines and early article sentences labeled as business, science/technology, sports, or world; (2) antonym, where each input word requires generating its opposite; (3) country-capital, comprising 197 country-capital pairs and requiring the capital city given a country; (4) country-currency, listing countries alongside their currencies; and (5) sentiment, which contains 1167 movie review sentences labeled as either positive or negative. Further, we translate (see Appendix for a discussion on high fidelity of the translations) these task datasets into specific languages and identify the attention heads.
Evaluation metrics: Following previous studies Chao et al. (2024); Qi et al. (2023b), we assess the safety standards of language models using the attack success rate (ASR). ASR measures the proportion of harmful responses generated by the model out of the total number of harmful queries and is defined as follows.
[…]
@@
GPT-4o as a safety evaluator: To reliably and efficiently assess a model’s safety performance, we employ GPT-4o as the annotator, leveraging its strong alignment with human judgments on harmful responses Qi et al. (2023b); Banerjee et al. (2024b); Hazra et al. (2024b). In particular, we adopt the evaluation prompt777see Appendix proposed by Banerjee et al. (2024b). Cases in which GPT-4o declines to provide annotations due to content filtering are excluded from our calculations. A lower ASR score indicates a safer model.